Systems Staff Meeting Notes 8/16/99

Systems Staff Meeting Notes
August 16, 1999

Present: Rich Bennett, Bill Covey, Winston Harris, Bobbie Meng, Ying Tang
  1. Analyst position - offer has been made, accepted. Hilton Soto - start date to be announced
  2. Net Coordinator - trying to pry resumes out of personnel at the Stadium
  3. Liaison training - doubled up on sessions on Friday, finished at 12:30
  1. Outage on Sunday - West, Education out. For information, should be interaction between us and CORE - call NERDC operations to page someone i.e. an engineer, no less than hour - call Bill and he will get someone.
  2. Moving West to Criser router, another heads up - possible building outage when that occurs.
  3. Two hour blip on Library Web pages on Saturday. Bill took remote control of the server and got it back up. Bill just noticed it, no one called about it.
  4. Great Flap at LAC - forbade Systems to work on the machines, then posted to list that Systems couldn't fix them. Tom Caswell and Leilani were dispatched - machines were misconfigured, did some tests to eliminate LAC's suspicion of a switch problem. Tom & LeiLani will try to diagnose the problem.
  1. Student projects are finished
  2. IR - enhancements added, waiting on feedback from Bobbie
  3. Mellon2 changes being worked on
  4. OPS tracking reports - Ying is working on them
  5. SQL7 server is quiet - will leave it alone, need to check on order
  6. Application software - Premier package of Office 2000 - Microsoft refused to sell to us, but will sell us Professional which has everything but Front Page & Photo Editor. FrontPage kick is $28.00/license, whole Office package is about $48.00
  1. Students back next week
  2. Scott is the only one here and only in the PM
  1. Debra out this week
  2. Suzy out next week

Back to the Systems Homepage

Last updated February 18, 2000
by Debra Harris